OpenStudy (anonymous):

help please! the force between two long parallel conductors is 15 kg/metre. the conductor spacing is 10 cm. If one conductor carries twice the current of the other, calculate the current in each conductor. help me analyze please.

OpenStudy (anonymous):

http://theory.uwinnipeg.ca/physics/mag/node10.html

OpenStudy (anonymous):

lilMissMindset: Don't ask the same question in multiple groups. "mathematics" is the wrong one anyways. -> http://openstudy.com/code-of-conduct

OpenStudy (shubhamsrg):

if you can tell, what is the expression of force per unit length between 2 infinite parallel current carrying wires ?

OpenStudy (anonymous):

@MuH4hA sorry. im going to erase my other questions on other groups.

OpenStudy (anonymous):

Okay :) (although - you _should_ ask physics-question in physics... but i am no mod, so idk - guess it's ok ;) )

OpenStudy (anonymous):

@shubhamsrg the expression of force per unit length b/w two conductors: (F2/L)=(u0*i1*i2)/2pi*d

OpenStudy (shubhamsrg):

@stgreen i knew that buddy and i wanted @lilMissMindset to ans my query, ofcorse//

OpenStudy (anonymous):

i see

OpenStudy (anonymous):

yea, saw it now. (F2/L)=(u0*i1*i2)/2pi*d

OpenStudy (shubhamsrg):

so where are you stuck..just plug in values given..cant be that difficult

OpenStudy (anonymous):

what value should i put in 2pi*D?

OpenStudy (anonymous):

is D the distance?

OpenStudy (shubhamsrg):

D is distance between the wires (in meters)

OpenStudy (anonymous):

how about permeability? what is its unit?

OpenStudy (anonymous):

oops sorry µ0 = 4π×10−7 ≈ 1.2566370614…×10−6 H·m−1 or N·A−2

OpenStudy (anonymous):

its henry per metre

OpenStudy (anonymous):

or newton per ampere?

OpenStudy (shubhamsrg):

dont worry about unit of Uo ,, just plug in everything in meters and kgs and you should get an ans in ampere..

OpenStudy (shubhamsrg):

all conventional units..

OpenStudy (anonymous):

aw, sorry, im being stupid. im having trouble with the unit. :(

OpenStudy (anonymous):

no no you u0 is in newton per ampere square..while the force you have is in Kg..convert your force in newtons first

OpenStudy (shubhamsrg):

i see..yes ,you have to convert that.. use mg = F for that conversion..

OpenStudy (anonymous):

aww. thank you so much! at last, i get this problem. :)

OpenStudy (anonymous):

@MuH4hA i got an answer in the mathematics group.

OpenStudy (anonymous):

MATHS section got a number of engineers too lady :P

OpenStudy (anonymous):

yea, that is why i put my questions in multiple groups, hoping for answers. not knowing, it is wrong according to codes of conduct. im one guilty student.

OpenStudy (anonymous):

uh-oh come on...its not such a CRIME lol

OpenStudy (anonymous):

right. lol

OpenStudy (anonymous):

thanks again. ;)